OpenStudy (anonymous):

curves: 1) x^2 + xy + y^2 = 1 2) x^2-xy+y^2=1 find dy/dx for each of the curves. can someone go with me point by point of how to solve it? i am stuck on xy and don't know what exactly its derivative, is it 1dy/dx or just 1.

OpenStudy (anonymous):

$$\frac{d}{dx}[xy]=y\frac{d}{dx}x+x\frac{d}{dx}y=y+xy'$$

OpenStudy (anonymous):

thank you, can you tell me how we got to this: \[1) \frac{ dy }{ dx } = \frac{ -2x+y }{ x+2y } \] \[2) \frac{ dy }{ dx } =\frac{ -2x+y }{ -x+2y }\]